Here's my question: Bob invested $20,000, part at 14% and part at 13%. If the total interest at the end of the year is $2,720, how much did he invest at 14%? I got $11,000, but I don't think that's right.
Answer by oscargut(2103)

C1 +C2 = 20000
C1(0,14)+C2(0,13)=2720
then
(0,14)C1+(0,14)C2=2800
-(0,14)C1-(0,13)C2=-2720
(0,01)C2=80 then C2=8000 and C1=12000